OpenStudy (anonymous):

Find the lowest common denominator that would eliminate the fractions in this equation. 4y/5-y/4+7/10=7/2

OpenStudy (agentnao):

20

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

4y/5-y/4+7/10=7/2 (20-> denominator) 8y-5y+14=70

OpenStudy (anonymous):

So the answer is 70 not 20?